This is a duplicate of Hall C log entry 171321 from 16:27:59 on October 26, 2008.
MPS BCM has been wrong by factor of 10. When the current at the injector at OLO2 is 2.0 uA ( when the faraday cup is inserted it agrees with this) the MPS BCM was reading 0.2 while the calibrated Hall C BCM was reading 1700nA. So the Hall C BCM and OLO2 was relative close. MCC check for beam losses and didn't find any. So the MPS BCM must be off by factor of 10.
MCC is calling to see what to do about the MPS BCM.